OpenStudy (anonymous):

Simplify the rational expression (3x-9)/(x^(2)-6x+9) and then find all #s that must be excluded from the domain of the simplified rational expression. I'm really just having trouble factoring it.

OpenStudy (zzr0ck3r):

\[\frac{3x-9}{x^2-6x+9}=\frac{3(x-3)}{(x-3)^2}=\frac{3}{x-3}\]

OpenStudy (zzr0ck3r):

@rrey understand?

OpenStudy (anonymous):

Can you explain to me why? I can sort of see it, but I need exact clarification.

OpenStudy (zzr0ck3r):

do you see why 3x-9 = 3(x-3)?

OpenStudy (zzr0ck3r):

x^2-6x+9 so we want two numbers that multiply to get 9 but add to get -6 -3 and -3 are two such numbers x^2-6x+9 = x^2-3x-3x+9 = x(x-3)-3(x-3) = (x-3)(x-3) = (x-3)^2

OpenStudy (anonymous):

well for the numerator, I see 3*x is 3x, and 3*-3 is -9 so im not trying to do anything with my positive 3, correct? i'm distributing it

OpenStudy (anonymous):

and yeah i udnerstand the bottom prt

OpenStudy (anonymous):

My textbook says it is \[\frac{ 3 }{ x-3 }\] am i missing something? @zzr0ck3r